OpenStudy (anonymous):

can someone help me with problems 9-11?

OpenStudy (anonymous):

9)a=20

OpenStudy (anonymous):

fist one start with \(2x+4=10\) then \[2x=6\] \[x=3\]

OpenStudy (anonymous):

10) a=18

OpenStudy (anonymous):

11) a=17

OpenStudy (anonymous):

\[3a+50=104\] \[3a=54\] \[a=18\] for number 10

OpenStudy (anonymous):

\[2a+30=4a-4\] \[30+4=4a-2a\] \[34=2a\] \[a=17\] for the last one